OpenStudy (anonymous):

The sum of two consecutive even integers is 226. What is the larger integer?

OpenStudy (anonymous):

assume the integer is x and the consecutive even integer is x+2 then x+x+2=226 2x=226-2 2x=224/2 x=112 the the larger integer is x+2 = 112+2 =114